Timothy owns 840 shares of Countess Corp., which is priced at $14.13 per share. The company plans a 3-for-5 reverse stock split. How many shares will Timothy own and what will the share price be after the reverse stock split?

504; $14.13

504; $8.48

504; $23.55

1,400; $8.48

1,400; $23.55
Answer #1

Stock will be reverse Split i.e. every 5 shares will be converted to 3 shares.

Value of Shares Tomithy have = No of Shares * Value per Share

= 840 * 14.13

= $ 11,869.2

So Timothy will have 840 * 3 /5 = 504 Shares after reverse Split

Value after reverse split remians same So, Share Price will be

$ 11,869.2 = New No of shares * Share Price

$ 11,869.2 = 504 * Share Price

Share Price = 11869.2 / 504 = $ 23.55

Option C is correct.

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.